### v1

# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains using test data. A highly effective CTTA method involves applying layer-wise adaptive learning rates, and selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. In this work, we aim to overcome these limitations by identifying layers through the quantification of model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity in order to approximate the domain shift, followed by adjusting their learning rates accordingly. Overall, this approach leads to a more robust and stable optimization than prior approaches. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C and demonstrate the efficacy of our method against standard benchmarks and prior methods.
1 Introduction
In a real-world setting, machine-perception systems [1] function in a rapidly changing environment. In such contexts, pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts. For instance, while these models may have been trained on clean images, images captured by sensors could be corrupted by various weather conditions.
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [6, 44]. Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [44]. While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation: since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s pre-trained knowledge due to continuous adaptations and large parameter updates across tasks.
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In [32, 8, 44, 37], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA [45] includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA [41] introduces the attachment of "meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
More recently, the authors in [21] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be "surgically" or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) [33] was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) [36] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
In this paper, we present a novel method called PALM to push the limit of adaptive learning rate methods for CTTA. Our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions. Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the KL divergence between the model output and a uniform distribution. This approach eliminates dependence on pseudo-labels and enables a more accurate estimation of layer-wise importance. Subsequently, the learning rates of the parameters of these layers are adjusted based on the degree of domain shift observed in the test data. We utilize parameter sensitivity as an indicator of the degree of distribution shift and employ weighted moving averages to aggregate sensitivity from previous data batches. Compared to existing adaptive learning rate methods, our proposed approach offers a better estimation of domain shift for adjusting learning rates.
To summarise, our contributions are as follows:
  • We present PALM, a novel approach to continual test-time adaptation. Specifically, we demonstrate improved schemes for adaptive learning rate methods for CTTA.
  • Considering the prediction uncertainty of the model at test-time, we design an approach to automatically select layers that would require further adaptation and push the learning rates of parameters of these layers for model updates, by estimating the domain shift based on parameter sensitivity.
  • Our method delivers promising results in both continual and gradual TTA across three standard benchmark datasets. In addition, we conduct and demonstrate ablation experiments on the different components of our work.
2 Related Works
Test-time adaptation (TTA): The overarching target of TTA is to adapt source models to unlabelled test data in real-time in a source-free online training fashion [16, 42, 44]. TENT [44] optimizes the Batch Normalisation (BN) parameters by minimizing the Shannon entropy [40] of the pseudo labels. In [37], the statistics of the BN parameters are replaced by the ones of the input corrupted data. Adacontrast [3] combines contrastive learning and utilizes the memory module to update the entire model’s parameters. While previous TTA methods showed significant improvements, they were demonstrated for a single domain shift at a time [4, 26, 37, 44] and did not consider continual distribution shifts.
Continual TTA (CTTA): CTTA poses an adaptation challenge wherein the model must continually adjust to evolving target data, accommodating a sequence of dynamic domain shifts. The first work, CoTTA [45], considers addressing this challenging task by setting up a teacher-student network, with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with forgetting and accumulation of errors with time. EcoTTA [41] introduces "meta networks" i.e. a BN layer and a convolution block are attached to the frozen source model. At every stage, an $L_{1}$ regularisation loss between the outputs of the meta network and that of the source model is computed to handle catastrophic forgetting.
Efficient fine-tuning approaches: Freezing model parameters to learn information from downstream data has witnessed a good body of work [39, 27, 51, 22, 7, 9]. Depending upon the distribution shift of the target task, in [21], the authors demonstrated that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive-learning rate methods: Regarding the studies on the importance of each layer in training, [47, 49] concluded that deeper layers extract task-specific features, as opposed to the shallow layers. Based on this finding, AutoLR [35] design an adapted learning rate (LR) scheme based on the weight variations between layers. Despite attempts to set different initial learning rates for each layer, these rates remain mostly constant throughout training. In the context of CTTA, LAW [33] used the FIM [36] to calculate the domain shift and automatically adjust the learning rates of parameters. However, to obtain the layer importance, the pseudo-labels are used to compute the FIM, which would be inaccurate. In addition, a direct accumulation of FIMs across domains would lead to error accumulation. Furthermore, the LR "weight" of each layer is upper-bounded by the base LR that is used, which ignores the relative "sensitivity" of parameters towards adaptation of the current batch and domain.
Parameter sensitivity: In the literature about neural network pruning, understanding the sensitivity of parameters towards training has been studied well [19, 28, 43, 29, 23]. The overall idea is that parameters with lower sensitivity need to be adapted more with an increased learning rate, and vice-versa. [23], in addition, introduces a temporal variation of sensitivity, due to uncertainty, as another factor during the training of models for natural language understanding and machine translation tasks.
3 Background
3.1 Problem Setting
The objective of continual test-time adaptation (CTTA) is to adapt a fixed discriminative pre-trained source model, parameterized by ${\theta_{s}}$ and trained on the source dataset ($\mathcal{X}^{source}$, $\mathcal{Y}^{source}$), to multiple incoming tasks $\mathcal{T}_{i}$ = {$x_{k}$}${}_{k=1}^{n}$ where $\mathcal{T}_{i}$ refers to the $i^{th}$ task and n is the total number of batches of $\mathcal{T}_{i}$, with changes in domain distributions between tasks. At each time step, ${\theta_{s}}$ is gradually adapted to each incoming batch ($x_{k}$) of task $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta}$. In this entire process of adaptation, the source dataset is unavailable for use, due to privacy and storage constraints. As evident, the source model is first parameterized by ${\theta_{s}}$ before the adaptation begins. With the continuum of tasks, at time step t, the complete model parameters ${\theta}^{t}$ consists of a matrix of parameters, given by ${\theta}^{t}$ = ${\theta^{t}_{1}}$, $\theta^{t}_{2}$,…, ${\theta^{t}_{n}}$,…., ${\theta^{t}_{N}}$}, where ${\theta^{t}_{n}}$, at time t, denotes the parameters of the $n^{th}$ trainable layer out of a total of N.
3.2 Analysis of Adaptive Learning Rate Schemes for CTTA
In this subsection, we provide a detailed view and analysis of surgical fine-tuning [21], AutoLR [35], and LAW [33]. This is essential to understand the design principles and reasons for having an adaptive learning rate scheme for CTTA, which also includes the motivation of our work.
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution, surgical fine-tuning [21] empirically show that heuristically fine-tuning only the first convolutional block results in a better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
Inspired by [47, 49], AutoLR [35] claims that the weight changes across layers should be in increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$, at epoch e, is defined as,
$\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$ (1)
They hypothesize that weight changes in the initial layers should be small because they extract general features whereas the deeper layers extract task-specific features. A constrained adaptive LR optimization scheme is set up to force the weight variations to increase with the layers without studying the importance of each layer during adaptation. However, the updated LRs, for each layer, do not deviate too much and remain constant throughout.
As the first adaptive LR work for CTTA, LAW [33] hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [21]. To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as,
$s(\theta_{n}^{t};x_{k})=\nabla_{\theta_{n}^{t}}\log(p_{\theta^{t}}(x_{k}))$ (2)
where $x_{k}$ and $\theta_{n}^{t}$ refer to the current test batch and parameters of layer n at time t. Then, the Fisher Information Matrix (FIM) [36], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. As it can be seen, the layer-wise importance (Eq. 2) completely relies on the pseudo-labels to compute the log-likelihood, which is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. Furthermore, the authors propose to capture "domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer during the adaptation, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as,
$\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$ (3)
Based on Eq. 3, this means that LAW [33] puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a "relatively" larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to an imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
Drawing from these primary observations, our method focuses on addressing the following: 1) Based on the prediction uncertainty of the model, we automatically select a subset of layers that would require further adaptation to a task, without relying on the pseudo-labels. 2) As an indicator of the domain shift, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates. We freeze the layers that are not selected.
4 Proposed Method
Figure 1: PALM framework: At time t, input batch $x_{k}$ is processed by the model, parameterized by $\theta^{t}$. The KL-divergence between the softmax predictions and a uniform distribution is backpropagated to select the layers with the gradient norm $\leq$ $\eta$, to quantify the uncertainty. The parameter sensitivities of these layers, as an indicator of domain shift meters, are computed to update their learning rates (Eq. 11). Finally, with the optimization objective in Eq. 16, we update the model with the adjusted learning rates of the parameters.
In this section, we introduce and formalize our proposed method for CTTA to address the key limitations of the works discussed above. Fig. 1 diagrammatically illustrates our entire approach.
4.1 Layer Selection by Quantifying Prediction Uncertainty
CTTA maximizes model performance in the current domain and is vital to assess prediction certainty amidst domain shifts. In the literature on continual learning and TTA, several works utilize FIM [18, 38, 33]. However, the unavailability of true labels, with an unknown distribution, motivates us to capture the prediction uncertainty. Prior literature proposed methods for uncertainty from the classifier output or feature space [12, 14, 24, 25, 20]. Motivated by [15], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence (KL) between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, $\frac{1}{C}$, …., $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C is the total number of classes. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, due to the distribution shift, the KL divergence should be smaller. Let h denote the classifier of the model parameterized by ${\theta^{t}}$. We compute the loss, denoting it as $\mathcal{L}(\theta^{t})$ at time stamp t, as follows,
$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(\hat{h}(x_{k}))||\text{{u}})$ (4)
where $\hat{h}(x_{k})$ is the classifier’s output logits ($h(x_{k})$) smoothed by a temperature coefficient T i.e. $\hat{h}(x_{k})$ = $h(x_{k})$/T. Our intuition is that batches within the same task are of the same distribution and so, smoothing the logits with large values of T will lead to them being indistinguishable. With model parameters $\theta^{t}$, the gradients (${{\nabla\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial{\theta}^{t}}$) can be simplified to,
${{\nabla\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{j=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),j)}{\partial\theta^{t}}$ (5)
On the whole, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq. 5). In test-time adaptation, since ground-truth labels are not present, the gradients are more indicative of familiarity of the model with the batch of data. For batches within the same task, that share the same distribution, the vector norm of Eq. 5 would be higher because of a higher KL divergence since the softmax predictions would be less uniformly distributed. However, if the norm is smaller, this would mean that the softmax outputs are closer to u. Our key idea is to select all the parameters of layers that have the norm of gradients to be less than or equal to a threshold $\eta$. Layers with small gradient norms would require more adaptation for the model to make a confident prediction. Since we freeze the remaining layers, this also helps to alleviate catastrophic forgetting to a large extent by preserving the prior knowledge. And so the scoring function, for each layer, is defined as,
$\mathcal{Z}_{\theta_{n}}=\|{{\nabla\mathcal{L}({\theta}_{n}^{t})}}\|_{1}$ (6)
As mentioned earlier, $\mathcal{Z}_{\theta_{n}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation. We freeze the layers with $\mathcal{Z}_{\theta_{n}}$ $>$ $\eta$ i.e. set their learning rate to 0. This way, in contrast to surgical fine-tuning [21], we automatically select a subset of layers for adaptation to a batch of data.
4.2 Parameter Sensitivity as an Indicator of Domain Shift
New studies have found that model parameters demonstrate levels of criticality [2, 50] across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [31, 34]. Inspired by this observation, we claim that parameters have different contributions during test-time adaptation. From Sec. 4.1, we adaptively select layers that are uncertain in their predictions. For parameters of these layers only, we compute their sensitivity i.e. an approximation of an error when it is removed.
In our work, motivated from [19, 30, 29], we compute the parameter sensitivity of a parameter $\theta_{n}^{t}$ based on the loss (Eq. 4) as follows,
$S_{n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{n}^{t}=0)$ (7)
Precisely, $S_{n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{n}^{t}$ is removed. Sensitivity, computed as a "weighting factor" of model weights [30, 29, 23], determines the level of adaptation needed for model parameters.
We utilize the gradients of parameters from the selected layers, computed by Eq. 5 to compactly derive Eq. 7 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ as,
$S_{n}^{t}=|\theta_{n}^{t}\nabla\mathcal{L}({\theta_{n}}^{t})|$ (8)
Our rationale is as follows: parameters with low sensitivity are pushed to have higher learning rates owing to more adaptation. However, $S_{n}^{t}$ of $\theta_{n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this "domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{n}^{t}$. This is formulated as,
$\hat{S}_{n}^{t}=\alpha S_{n}^{t}+(1-\alpha)\hat{S}_{n}^{(t-1)}$ (9)
where $\hat{S}_{n}^{t}$ denotes the "domain-level" sensitivity and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{n}^{t}$ to $\hat{S}_{n}^{t}$. Since CTTA happens at batch level, it is critical to achieve a good balance between $S_{n}^{t}$ and $\hat{S}_{n}^{t}$, with a slightly larger focus on $S_{n}^{t}$. As it can be seen, we do not directly accumulate the "domain-level" information, unlike LAW [33].
In addition, due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. To quantify these statistics, an absolute difference between $\hat{S}_{n}^{t}$ and $S_{n}^{t}$ can be computed as,
$\hat{D}_{n}^{t}=|S_{n}^{t}-\hat{S}_{n}^{t}|$ (10)
We emphasize that since Eq. 9 leverages sensitivity from previously computed batches and domains, $\hat{D}_{n}^{t}$ would act as a good indicator of how uncertain $S_{n}^{t}$ is to the previous batches and domains, for $\theta_{n}$. This would take care of error accumulation too.
To automatically alter the learning rates of these parameters, the learning rate "importance" ($i_{n}^{t}$) is calculated as follows,
$i_{n}^{t}=(\hat{D}_{n}^{t})/(\hat{S}_{n}^{t})$ (11)
Then, the final learning rate to update $\theta_{n}^{t}$ would be
$\hat{\kappa}_{n}^{t}=\kappa*i_{n}^{t}$ (12)
where $\kappa$ is the base fixed learning rate. The intuition behind the formulation of Eq. 11 is as follows - smaller values of $\hat{S}_{n}^{t}$ indicate lower sensitivity and hence larger parameter updates are required via promoting the learning rate. At the same time, large uncertainties by $\hat{D}_{n}^{t}$ would also require larger updates. To avoid any overflow/underflow, we add $\epsilon$ (0<$\epsilon$$\ll$1) to the numerator and the denominator in Eq. 11.
4.3 Optimization Objective
Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [44] minimized the Shannon entropy on the output logits, defined as,
$\mathcal{H}(\hat{y})=-\sum_{c}p(y_{c})log(p(y_{c}))$ (13)
where $p(y_{c})$ refers to the softmax probability of class c i.e. $p(y_{c})$ = softmax($h(x_{k})$). Following previous CTTA work [41], we use entropy minimization as the main loss objective for adaptation where Eq. 13 is minimized for samples below a certain threshold.
$\mathcal{L}^{entropy}=\mathds{1}_{\mathcal{H}(\hat{y})\leq\mathcal{H}_{0}}\mathcal{H}(\hat{y})$ (14)
where $\mathds{1}$ is the indicator function and $\mathcal{H}_{0}$ is a fixed threshold set to $0.4*lnC$, consistently following [41]. To be fair with prior works [45, 33], we use the consistency loss between the model predictions on $x_{k}$ and its augmented version $\hat{x}_{k}$ to regularize the model as follows,
$\mathcal{L}^{const}=-\sum_{c}p(y_{c})log(p(\hat{y}_{c}))$ (15)
where, $\hat{y}_{c}$ = $h(\hat{x}_{k})$. Our overall optimization objective is as follows,
$\mathcal{L}^{total}=\mathcal{L}^{entropy}+\lambda\mathcal{L}^{const}$ (16)
where $\lambda$ indicates a regularisation coefficient to balance the consistency loss. In summary, upon calculating the parameter-specific learning rates (Eq. 12), we optimize the continually fine-tuned model to minimize Eq. 16.
5 Experiments
We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting. However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [45, 33]. In the following subsections, we describe the implementation details of both the challenging adaptation settings. For additional results and analysis, please refer to the Supplementary Material.
5.1 Experimental Settings and Baselines
Datasets and tasks. Following the standard benchmarks in CTTA, set by CoTTA [45], we evaluate our proposed method on datasets - CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [11]. Each dataset contains a set of 15 corruption styles (e.g. gaussian noise, shot noise, impulse noise, defocus blur) with 5 severity levels. In line with established practices outlined in [45, 6], we report the mean classification error as the evaluation metric.
CTTA implementation details. We evaluate our approach using frameworks from prior state-of-the-art methods [44, 45, 3, 41, 37, 33]. To maintain fairness regarding the usage of pre-trained vision models, we employ WideResNet-28 [48] for CIFAR-10C, ResNeXt-29 [46] for CIFAR-100C, and ResNet-50 [10] for ImageNet-C, all available on RobustBench [5]. We optimize using an Adam optimizer [17], setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA [45], and results are reported for the highest severity level of 5 for each task.
GTTA implementation details. In the standard CTTA setup, each task receives images with the highest severity. However, in GTTA, following [45, 41, 33], for each task the severity levels are gradually changed as …3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3… We maintain consistency with the experimental details outlined in the CTTA setting.
Baselines. We compare our work against several CTTA baselines, including Source, BN Stats Adapt [37], TENT Continual [44], CoTTA [45], EcoTTA [41], Adacontrast [3], and LAW [33]. In addition, we also contrast our work against surgical fine-tuning [21]. For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [46] for CIFAR-100C. Following the findings of surgical fine-tuning [21], we continually fine-tune only the first convolutional block of the respective models, denoted as Surgical fine-tuning (continual). We have mentioned the training details in the Supplementary. For the reported results of LAW [33], we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
Table 1: Mean errors (in %) on CIFAR-10C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using WideResNet-28 [48] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
BN Stats Adapt [37] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT Continual [44] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA [45]
24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
ECoTTA [41]
23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
Adacontrast [3]
29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
Surgical
fine-tuning [21]
27.71 24.14 33.79 12.18 33.07 13.09 10.97 15.94 15.61 13.8 7.84 10.77 21.17 16.62 22.88 18.64
LAW [33]
24.71 18.91 25.56 12.96 26.78 15.09 11.8 15.17 14.78 15.89 10.13 13.89 19.41 14.74 18.35 17.21
Ours
25.89 18.1 22.71 12.38 25.3 13.18 10.79 13.5 13.15 12.21 8.53 11.85 17.98 12.02 15.47 15.54
Table 2: Mean errors (in %) on CIFAR-100C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNeXt-29 [46] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source
73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN Stats Adapt [37]
42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
TENT Continual [44]
37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
CoTTA [45]
40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
Adacontrast [3]
42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
Surgical
fine-tuning [21]
40.75 36.36 37.89 25.53 39.02 27.65 25.27 32.26 30.45 37.35 23.78 26.82 32.37 27.64 37.20 32.02
LAW [33]
41.06 36.73 38.38 25.67 37.02 27.82 25.22 30.79 30.04 37.35 24.41 27.69 31.25 27.84 34.92 31.75
Ours
37.33 32.5 34.94 26.20 35.28 27.54 24.64 28.85 29.15 34.14 23.52 26.99 31.15 26.56 34.1 30.19
5.2 CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C. We contrast our proposed method against several baselines including surgical fine-tuning [21]. Tables 1 and 2 report the CTTA mean classification errors across all the domains, for each of the methods. The results reveal that our method performs the best across all training setups and baselines for both datasets. In particular, we achieve large improvements of 4.86% and 5.16% for CIFAR-10C and 5.21% and 30.71% for CIFAR-100C, over BN Stats Adapt[3] and TENT Continual[44] respectively. Concerning recent work LAW [33] that uses an adaptive LR scheme, our method achieves an improvement of 1.67% and 1.56% on CIFAR-10C and CIFAR-100C, respectively. It’s noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation over time, leading to an increase in mean error. In the study by CoTTA [45], despite achieving good performance, the approach’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. LAW [33] uses the uncertain pseudo labels, of each batch, to calculate the domain shift approximated via the FIM computation. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the "domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In comparison, the computation of the model prediction uncertainty to select layers is useful in our work (Eq. 4). Another interesting observation is that with a drastic reduction in the number of trainable parameters, as in Surgical fine-tuning (continual), the mean errors in Table 2 is better than most CTTA benchmarks, if not the best. This affirms the findings in [21] that concentrated adaptation, for image corruptions, in the initial layers is sufficient.
Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset, with ResNet-50 pre-trained on ImageNet. As shown, our method outperforms all the previous CTTA baselines and Surgical fine-tuning (continual). We observe an improvement of 1% over LAW [33]. In addition, we achieve 2.31%, 5.38%, and 30.71% improvements over CoTTA [45], Adacontrast [3], and TENT [44], respectively.
Table 3: Mean errors (in %) on ImageNet-C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNet-50 [10] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source
97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
BN Stats Adapt [37]
85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
TENT Continual [44]
81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
CoTTA [45]
84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
ECoTTA [41]
- - - - - - - - - - - - - - - 63.4
Adacontrast [3]
82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
Surgical
fine-tuning [21]
83.42 79.5 77.28 86.1 82.4 72.28 61.18 63.1 64.76 51.26 35.92 72.9 56.24 46.72 50.2 65.55
LAW [33]
81.86 75.04 72.12 77.42 73.22 63.96 54.6 57.92 61.06 50.1 36.36 68.66 49.08 46.22 49.12 61.12
Ours
81.06 73.32 70.88 77.00 71.88 62.34 53.94 56.7 60.76 50.36 36.26 65.96 48.08 45.28 48.0 60.12
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the gradual TTA (GTTA) mean errors averaged across all the severity levels (1normal-$\rightarrow$2normal-$\rightarrow$3normal-$\rightarrow$4normal-$\rightarrow$5).
Dataset Source
BN Stats
Adapt
TENT
Continual
CoTTA Adacontrast
Surgical
fine-tuning
LAW Ours
CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.33 10.87 9.56
CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.31 27.19 26.11
ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.36 40.52 38.49
5.3 GTTA Classification Results
Results on CIFAR-10C, CIFAR-100C, and Imagenet-C. This subsection presents the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C. For each task, the corruption severity of the inputs is gradually increased from levels 1 to 5, as discussed in the implementation details in Sec. 5.1. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the baselines [37, 44, 45, 3, 33]. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA [45], which also demonstrates competitive improvements in the mean classification error.
5.4 Ablation Studies
We conduct ablation experiments to validate our approach and gain insights into its underlying mechanisms and components. Specifically, we conduct these experiments and report results on the three CTTA benchmark datasets: CIFAR-10C, CIFAR-100C, and ImageNet-100C.
    WideResNet-28 [48]       ResNeXt-29 [46]       ResNet-50 [10]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Illustrations of the learning rate importances of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10] during CTTA, for different convolutional blocks/stages/layers. [Top row] - Variations across respective datasets for domain glass_blur. [Bottom row] - Variations across respective datasets for domain snow.
Analysis of the learning rate importance. The prime objective of our work is to adaptively alter the learning rates of parameters of selected layers based on their sensitivity. In Fig. 2, we illustrate how the learning rate (LR) importance (Eq. 11) of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10], averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. As we can see, the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated LR importances, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning [21]. In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Nullifying prediction uncertainty. In LAW [33], the authors make use of the pseudo-labels to automatically decide on the importance of layers for adaptation. In this, we study the effectiveness of such a design for our method i.e. we investigate if layer importance computed using pseudo-labels is helpful. We compute the gradients from the log-likelihood (Eq. 2) and estimate the sensitivity of the parameters of the chosen layers. Table 5.4 reports the mean classification errors for all the datasets. We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty. The scoring function (Eq. 6) computed using these gradients carries insufficient information. We achieve improvements of 2.44%, 3.69%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
Refer to caption Refer to caption
a) Ablation on smoothing factor $\alpha$. b) Ablation on temperature T.
Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T for CIFAR-10C, CIFAR-100C, and ImageNet-C.
Table 5: Prediction uncertainty using pseudo-labels vs our method. Dataset Pseudo-labels Ours CIFAR-10C 17.98 15.54 CIFAR-100C 33.88 30.19 ImageNet-C 68.62 60.12
Table 6: Ablation results on benchmark datasets with varying $\lambda$. $\lambda$ 0 0.01 0.05 0.1 0.5 1.0 CIFAR-10C 16.84 15.54 16.45 17.03 17.65 17.93 CIFAR-100C 30.44 30.19 30.51 30.57 30.63 30.7 ImageNet-C 63.98 60.12 60.38 60.4 60.4 60.42
Contribution of smoothing factor $\alpha$. An important component of our method is the computation of the moving average of sensitivities across data batches, smoothed by a parameter $\alpha$ (Eq. 9). Fig. 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the "domain-level" sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e. when the current sensitivity of parameters of selected layers is not considered, the mean classification error for the respective datasets drastically spikes up. This suggests that the effect of the current sensitivity of selected layers is indeed required. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.67% over LAW [33]. Similarly, for CIFAR-100C and ImageNet-C, we see improvements of 1.56% and 1% over LAW [33], at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much and are stable. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{n}^{t}$) and current sensitivity ($S_{n}^{t}$) is critical, as seen in the figure.
[…]
Robustness to the regularisation coefficient $\lambda$. As an important ablation step, we understand the effect of the parameter $\lambda$ on the final optimization objective, as in Eq. 16. We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. We study the contributions of the consistency loss to entropy minimization. Table 5.4 reports the average classification errors on CIFAR-10C, CIFAR-100C, and ImageNet-C, with varying $\lambda$. For $\lambda$ = 0.01, we achieve the best results across all the datasets. In addition, we observe that our method is very robust to $\lambda$ i.e. with subtle variations to differing $\lambda$, across the datasets. In particular, for CIFAR-10C, with $\lambda$ = 0 i.e. with no consistency loss, we achieve improvements of 0.37% and 3.86% over LAW [33] and TENT Continual [44]. For CIFAR-100C, with no consistency loss ($\lambda$ = 0), we notice improvements of 1.31%, 2.06%, 4.96%, and 30.46% over LAW [33], CoTTA [45], BN Stats Adapt [37] and TENT Continual [44]. As for ImageNet-C, the consistent best setting of $\lambda$ = 0.01 is better than LAW [33] by 1% and by 5.38% over Adacontrast [3]. However, with $\lambda$ = 0 i.e. no consistency loss, we see a spike in the error rate by about 3.86% over the best setting ($\lambda$ = 0.01).
6 Conclusion
In this paper, we propose a novel approach called PALM, to continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e. by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.